# Anisotropy of the anomalous Hall effect in the altermagnet candidate MnSi films

### v1

@@ III Results and discussions
Although the exact Néel vector orientation and rotation under the action of the external magnetic field cannot be deduced from the results presented, in this final paragraph we wish to open a discussion on the interpretation of the anisotropic AHE in the light of the symmetries of the epitaxial MnSi considering this material to be a strong candidate for the altermagnetic phase. First, it should be noted that in our measurement geometry (Figure 1), the changes in correspond to changes in the component of along the [0001] direction (out-of-plane). In altermagnets, the presence and orientation of depend on the -vector orientation because the latter influences the symmetries of the system. While accurate calculation of the entire -dependence of is inhibited by the unknown relevant micromagnetic parameters, the value of for some specific orientations can be inferred from symmetry-considerations. The symmetry analysis on the expected spin structure of the epitaxial MnSi shown in Figure 1(a) indicates that for non-zero the -vector cannot point along the [0001] axis ( excluded), along the main in-plane axes (only in-plane allowed), or in the planes of the high-symmetry in-plane axes highlighted in Figure 1(a) (only in-plane allowed). Here, the crystal symmetries of the planes corresponding to 0, 60, and 120 deg and 30, 90, and 150 deg differ, as seen in Figure 5 and therefore the likely differ, too.
@@ IV Conclusion
In conclusion, the implications of our measurements on the external field orientation dependence of in epitaxial MnSi films are two-fold. First, the systematic changes in supported by the systematic changes in AMR upon the field rotation underlines our ability to manipulate the -orientation with an external field of the order of 2 T or more. This is an important premise for further experimentally feasible spin transport studies in altermagnets where control over the -orientation is critical. Second, our key result is that the in epitaxial MnSi films shows anisotropic behavior as a function of the external field-induced orientation of relative to the underlying crystal structure, and this anisotropic behavior depends on the crystal quality and is unlikely to relate to the magnetocrystalline anisotropy. This observation of the anisotropic adds to the growing array of properties of epitaxial MnSi advocating the altermagnetic phase in this material.
@@ Acknowledgements.
This work was supported by the French national research agency (ANR) and the Deutsche Forschungsgemeinschaft (DFG) (Project MATHEEIAS- Grant No. ANR-20-CE92-0049-01 / DFG-445976410). H. R., E. S., S. T. B. G. and A. T. are supported by the DFG-GACR grant no. 490730630. J. R. acknowledges MINECO for the Margaritas Salas program. D. K. acknowledges the academy of sciences of the Czech Republic for the Lumina quaeruntur program.References
@@
- Nayak et al. [2016] A. K. Nayak, J. E. Fischer, Y. Sun, B. Yan, J. Karel, A. C. Komarek, C. Shekhar, N. Kumar, W. Schnelle, J. Kübler, C. Felser, and S. S. P. Parkin, Large anomalous Hall effect driven by a nonvanishing Berry curvature in the noncolinear antiferromagnet MGe, Science Advances 2, e1501870 (2016).
@@
- Tschirner et al. [2023] T. Tschirner, P. Keßler, R. D. Gonzalez Betancourt, T. Kotte, D. Kriegner, B. Büchner, J. Dufouleur, M. Kamp, V. Jovic, L. Smejkal, J. Sinova, R. Claessen, T. Jungwirth, S. Moser, H. Reichlova, and L. Veyrat, Saturation of the anomalous hall effect at high magnetic fields in altermagnetic ruo2, APL Materials 11, 10.1063/5.0160335 (2023).
@@
- Reichlová et al. [2021] H. Reichlová, R. L. Seeger, R. González-Hernández, I. Kounta, R. Schlitz, D. Kriegner, P. Ritzinger, M. Lammel, M. Leiviskä, V. Petříček, P. Doležal, E. Schmoranzerová, A. Bad’ura, A. Thomas, V. Baltz, L. Michez, J. Sinova, S. T. B. Goennenwein, T. Jungwirth, and L. Šmejkal, Macroscopic time reversal symmetry breaking by staggered spin-momentum interaction (2021), arXiv:2012.15651 [cond-mat.mes-hall] .
- Kounta et al. [2023] I. Kounta, H. Reichlova, D. Kriegner, R. Lopes Seeger, A. Bad’ura, M. Leiviska, A. Boussadi, V. Heresanu, S. Bertaina, M. Petit, E. Schmoranzerova, L. Smejkal, J. Sinova, T. Jungwirth, V. Baltz, S. T. B. Goennenwein, and L. Michez, Competitive actions of MnSi in the epitaxial growth of MnSi thin films on Si(111), Phys. Rev. Mater. 7, 024416 (2023).
@@
- Fedchenko et al. [2023] O. Fedchenko, J. Minar, A. Akashdeep, S. W. D’Souza, D. Vasilyev, O. Tkach, L. Odenbreit, Q. L. Nguyen, D. Kutnyakhov, N. Wind, L. Wenthaus, M. Scholz, K. Rossnagel, M. Hoesch, M. Aeschlimann, B. Stadtmueller, M. Klaeui, G. Schoenhense, G. Jakob, T. Jungwirth, L. Smejkal, J. Sinova, and H. J. Elmers, Observation of time-reversal symmetry breaking in the band structure of altermagnetic RuO (2023), arXiv:2306.02170 .
@@
- Lee et al. [2023] S. Lee, S. Lee, S. Jung, J. Jung, D. Kim, Y. Lee, B. Seok, J. Kim, B. G. Park, L. Šmejkal, C.-J. Kang, and C. Kim, Broken Kramers’ degeneracy in altermagnetic MnTe (2023), arXiv:2308.11180 .
@@
- Osumi et al. [2023] T. Osumi, S. Souma, T. Aoyama, K. Yamauchi, A. Honma, K. Nakayama, T. Takahashi, K. Ohgushi, and T. Sato, Observation of giant band splitting in altermagnetic MnTe (2023), arXiv:2308.10117 .
@@
- Reimers et al. [2023] S. Reimers, L. Odenbreit, L. Smejkal, V. N. Strocov, P. Constantinou, A. B. Hellenes, R. J. Ubiergo, W. H. Campos, V. K. Bharadwaj, A. Chakraborty, T. Denneulin, W. Shi, R. E. Dunin-Borkowski, S. Das, M. Kläui, J. Sinova, and M. Jourdan, Direct observation of altermagnetic band splitting in CrSb thin films (2023), arXiv:2310.17280 .
@@
- Sattigeri et al. [2023] R. M. Sattigeri, G. Cuono, and C. Autieri, Altermagnetic surface states: towards the observation and utilization of altermagnetism in thin films, interfaces and topological materials, Nanoscale 15, 16998–17005 (2023).
@@
- Zhou et al. [2023] X. Zhou, W. Feng, R.-W. Zhang, L. Smejkal, J. Sinova, Y. Mokrousov, and Y. Yao, Crystal thermal transport in altermagnetic RuO (2023), arXiv:2305.01410 .
- González-Hernández et al. [2021] R. González-Hernández, L. Šmejkal, K. Výborný, Y. Yahagi, J. Sinova, T. c. v. Jungwirth, and J. Železný, Efficient electrical spin splitter based on nonrelativistic collinear antiferromagnetism, Phys. Rev. Lett. 126, 127701 (2021).
@@
- Bose et al. [2022] A. Bose, N. J. Schreiber, R. Jain, D.-F. Shao, H. P. Nair, J. Sun, X. S. Zhang, D. A. Muller, E. Y. Tsymbal, D. G. Schlom, and D. C. Ralph, Tilted spin current generated by the collinear antiferromagnet ruthenium dioxide, Nature Electronics 5, 267–274 (2022).
@@
- Bai et al. [2022] H. Bai, L. Han, X. Y. Feng, Y. J. Zhou, R. X. Su, Q. Wang, L. Y. Liao, W. X. Zhu, X. Z. Chen, F. Pan, X. L. Fan, and C. Song, Observation of spin splitting torque in a collinear antiferromagnet RuO, Phys. Rev. Lett. 128, 197202 (2022).
@@
- Chi et al. [2023] B. Chi, L. Jiang, Y. Zhu, G. Yu, C. Wan, J. Zhang, and X. Han, Crystal facet orientated altermagnets for detecting ferromagnetic and antiferromagnetic states by giant tunneling magnetoresistance effect (2023), arXiv:2309.09561 .
@@
- Lander et al. [1967] G. H. Lander, P. J. Brown, and J. B. Forsyth, The antiferromagnetic structure of MnSi, Proceedings of the Physical Society 91, 332 (1967).
- Menshikov et al. [1990] A. Z. Menshikov, A. P. Vokhmyanin, and Y. A. Dorofeev, Magnetic structure and phase transformations in MnSi, physica status solidi (b) 158, 319 (1990).
- Brown et al. [1992] P. J. Brown, J. B. Forsyth, V. Nunez, and F. Tasset, The low-temperature antiferromagnetic structure of MnSi revised in the light of neutron polarimetry, Journal of Physics: Condensed Matter 4, 10025 (1992).
- Brown and Forsyth [1995] P. J. Brown and J. B. Forsyth, Antiferromagnetism in MnSi: the magnetic structure of the AF2 phase at 70 K, Journal of Physics: Condensed Matter 7, 7619 (1995).

### v2

@@ III Results and discussions
Although the exact Néel vector orientation and rotation under the action of the external magnetic field cannot be deduced from the results presented, we wish to open a discussion on the interpretation of the anisotropic AHE in the light of the symmetries of the epitaxial Mn5Si3. This material is considered to be a candidate for the altermagnetic phase [17, 18, 19, 28, 29]. First, it should be noted that in our measurement geometry (Figure 1), the changes in correspond to changes in the component of along the [0001] direction (out-of-plane). In altermagnets, the presence and orientation of depend on the -vector orientation because the latter influences the symmetries of the system. While accurate calculation of the entire -dependence of is inhibited by the unknown relevant micromagnetic parameters, the value of for some specific orientations can be inferred from symmetry considerations. More specifically, considering the expected spin structure of the epitaxial Mn5Si3 shown in Figure 1(a) we can use symmetry analysis to obtain information on the -vector orientations that give zero in the xy direction: if -vector is along [0001] crystal axis, is forbidden in any direction due to , , and symmetries, while if -vector is along any high-symmetry in-plane axis (highlighted in Figure 1(a)), only in-plane is allowed due to the and symmetries. This illustrates the influence of the orientation on the through symmetry, and thus the fact that the symmetries of the planes corresponding to 0, 60, and 120 deg and 30, 90, and 150 deg are different (Figure 5) can give a plausible explanation to the anisotropy of in these planes.
@@ IV Conclusion
In conclusion, the implications of our measurements on the external field orientation dependence of in epitaxial Mn5Si3 films are two-fold. First, the systematic changes in supported by the systematic changes in AMR|| upon the field rotation underlines our ability to manipulate the -orientation with an external field of the order of 2 T and above. This is an important premise for further experimentally feasible spin transport studies in altermagnets where control over the -orientation is critical. Second, our key result is that the in epitaxial Mn5Si3 films shows anisotropic behavior as a function of the external field-induced orientation of relative to the underlying crystal structure, and this anisotropic behavior depends on the crystal quality and is unlikely to relate to the magnetocrystalline anisotropy. This observation of the anisotropic adds to the growing array of properties of epitaxial Mn5Si3: theoretical prediction of the collinear compensated hexagonal phase required and experimental demonstrations of vanishing net magnetization, strong spontaneous anomalous Hall [17, 19] and Nernst [28, 29] effects , and their dependence on the crystallinity- [18] and Mn-content[29]. These properties together advocate the altermagnetic phase in this material.
@@ Acknowledgements.
This work was supported by the French national research agency (ANR) and the Deutsche Forschungsgemeinschaft (DFG) (Project MATHEEIAS- Grant No. ANR-20-CE92-0049-01 / DFG-445976410). H. R., E. S., S. T. B. G. and A. T. are supported by the DFG-GACR grant no. 490730630. J. R. acknowledges MINECO for the Margaritas Salas program. D. K. acknowledges the academy of sciences of the Czech Republic for the Lumina quaeruntur program. AB,ES,DK, LS,TJ acknowledge the Ministry of Education, Youth and Sports of the Czech Republic through the OP JAK call Excellent Research (TERAFIT Project No. CZ.02.01.01/00/22_008/0004594)References
@@
- Nayak et al. [2016] A. K. Nayak, J. E. Fischer, Y. Sun, B. Yan, J. Karel, A. C. Komarek, C. Shekhar, N. Kumar, W. Schnelle, J. Kübler, C. Felser, and S. S. P. Parkin, Large anomalous Hall effect driven by a nonvanishing Berry curvature in the noncolinear antiferromagnet M3Ge, Science Advances 2, e1501870 (2016).
@@
- Hayami et al. [2019] S. Hayami, Y. Yanagi, and H. Kusunose, Momentum-dependent spin splitting by collinear antiferromagnetic ordering, Journal of the Physical Society of Japan 88, 123702 (2019).
- Yuan et al. [2020] L.-D. Yuan, Z. Wang, J.-W. Luo, E. I. Rashba, and A. Zunger, Giant momentum-dependent spin splitting in centrosymmetric low- antiferromagnets, Phys. Rev. B 102, 014422 (2020).
@@
- Tschirner et al. [2023] T. Tschirner, P. Keßler, R. D. Gonzalez Betancourt, T. Kotte, D. Kriegner, B. Büchner, J. Dufouleur, M. Kamp, V. Jovic, L. Smejkal, J. Sinova, R. Claessen, T. Jungwirth, S. Moser, H. Reichlova, and L. Veyrat, Saturation of the anomalous hall effect at high magnetic fields in altermagnetic ruo2, APL Materials 11, 101103 (2023).
@@
- Reichlová et al. [2021] H. Reichlová, R. L. Seeger, R. González-Hernández, I. Kounta, R. Schlitz, D. Kriegner, P. Ritzinger, M. Lammel, M. Leiviskä, V. Petříček, P. Doležal, E. Schmoranzerová, A. Bad’ura, A. Thomas, V. Baltz, L. Michez, J. Sinova, S. T. B. Goennenwein, T. Jungwirth, and L. Šmejkal, Macroscopic time reversal symmetry breaking by staggered spin-momentum interaction (2021), arXiv:2012.15651 .
- Kounta et al. [2023] I. Kounta, H. Reichlova, D. Kriegner, R. Lopes Seeger, A. Bad’ura, M. Leiviska, A. Boussadi, V. Heresanu, S. Bertaina, M. Petit, E. Schmoranzerova, L. Smejkal, J. Sinova, T. Jungwirth, V. Baltz, S. T. B. Goennenwein, and L. Michez, Competitive actions of MnSi in the epitaxial growth of Mn5Si3 thin films on Si(111), Phys. Rev. Mater. 7, 024416 (2023).
- Han et al. [2024a] L. Han, X. Fu, R. Peng, X. Cheng, J. Dai, L. Liu, Y. Li, Y. Zhang, W. Zhu, H. Bai, Y. Zhou, S. Liang, C. Chen, Q. Wang, X. Chen, L. Yang, Y. Zhang, C. Song, J. Liu, and F. Pan, Electrical 180° switching of néel vector in spin-splitting antiferromagnet, Science Advances 10, eadn0479 (2024a).
@@
- Mazin et al. [2021] I. I. Mazin, K. Koepernik, M. D. Johannes, R. González-Hernández, and L. Šmejkal, Prediction of unconventional magnetism in doped fesb 2, Proceedings of the National Academy of Sciences 118, 10.1073/pnas.2108924118 (2021).
@@
- Fedchenko et al. [2024] O. Fedchenko, J. Minar, A. Akashdeep, S. W. D’Souza, D. Vasilyev, O. Tkach, L. Odenbreit, Q. L. Nguyen, D. Kutnyakhov, N. Wind, L. Wenthaus, M. Scholz, K. Rossnagel, M. Hoesch, M. Aeschlimann, B. Stadtmueller, M. Klaeui, G. Schoenhense, G. Jakob, T. Jungwirth, L. Smejkal, J. Sinova, and H. J. Elmers, Observation of time-reversal symmetry breaking in the band structure of altermagnetic RuO2, Science Advances 10, eadj4883 (2024).
@@
- Lee et al. [2024] S. Lee, S. Lee, S. Jung, J. Jung, D. Kim, Y. Lee, B. Seok, J. Kim, B. G. Park, L. Šmejkal, C.-J. Kang, and C. Kim, Broken kramers degeneracy in altermagnetic mnte, Phys. Rev. Lett. 132, 036702 (2024).
@@
- Osumi et al. [2024] T. Osumi, S. Souma, T. Aoyama, K. Yamauchi, A. Honma, K. Nakayama, T. Takahashi, K. Ohgushi, and T. Sato, Observation of a giant band splitting in altermagnetic mnte, Phys. Rev. B 109, 115102 (2024).
@@
- Reimers et al. [2024] S. Reimers, L. Odenbreit, L. Smejkal, V. N. Strocov, P. Constantinou, A. B. Hellenes, R. J. Ubiergo, W. H. Campos, V. K. Bharadwaj, A. Chakraborty, T. Denneulin, W. Shi, R. E. Dunin-Borkowski, S. Das, M. Kläui, J. Sinova, and M. Jourdan, Direct observation of altermagnetic band splitting in CrSb thin films, Nature Communications 15, 2116 (2024).
@@
- Sattigeri et al. [2023] R. M. Sattigeri, G. Cuono, and C. Autieri, Altermagnetic surface states: towards the observation and utilization of altermagnetism in thin films, interfaces and topological materials, Nanoscale 15, 16998 (2023).
@@
- Zhou et al. [2024] X. Zhou, W. Feng, R.-W. Zhang, L. Šmejkal, J. Sinova, Y. Mokrousov, and Y. Yao, Crystal thermal transport in altermagnetic , Phys. Rev. Lett. 132, 056701 (2024).
- Badura et al. [2024] A. Badura, W. H. Campos, V. K. Bharadwaj, I. Kounta, L. Michez, M. Petit, J. Rial, M. Leiviskä, V. Baltz, F. Krizek, D. Kriegner, J. Zemen, S. Telkamp, S. Sailler, M. Lammel, R. J. Ubiergo, A. B. Hellenes, R. González-Hernández, J. Sinova, T. Jungwirth, S. T. B. Goennenwein, L. Šmejkal, and H. Reichlova, Observation of the anomalous nernst effect in altermagnetic candidate mn5si3 (2024), arXiv:2403.12929 .
- Han et al. [2024b] L. Han, X. Fu, W. He, Y. Zhu, J. Dai, W. Yang, W. Zhu, H. Bai, C. Chen, C. Wan, X. Han, C. Song, J. Liu, and F. Pan, Observation of non-volatile anomalous nernst effect in altermagnet with collinear néel vector (2024b), arXiv:2403.13427 .
- Naka et al. [2019] M. Naka, S. Hayami, H. Kusunose, Y. Yanagi, Y. Motome, and H. Seo, Spin current generation in organic antiferromagnets, Nature Communications 10, 4305 (2019).
- Ma et al. [2021] H.-Y. Ma, M. Hu, N. Li, J. Liu, W. Yao, J.-F. Jia, and J. Liu, Multifunctional antiferromagnetic materials with giant piezomagnetism and noncollinear spin current, Nature Communications 12, 4305 (2021).
@@
- Bose et al. [2022] A. Bose, N. J. Schreiber, R. Jain, D.-F. Shao, H. P. Nair, J. Sun, X. S. Zhang, D. A. Muller, E. Y. Tsymbal, D. G. Schlom, and D. C. Ralph, Tilted spin current generated by the collinear antiferromagnet ruthenium dioxide, Nature Electronics 5, 267 (2022).
@@
- Bai et al. [2022] H. Bai, L. Han, X. Y. Feng, Y. J. Zhou, R. X. Su, Q. Wang, L. Y. Liao, W. X. Zhu, X. Z. Chen, F. Pan, X. L. Fan, and C. Song, Observation of spin splitting torque in a collinear antiferromagnet RuO2, Phys. Rev. Lett. 128, 197202 (2022).
- Karube et al. [2022] S. Karube, T. Tanaka, D. Sugawara, N. Kadoguchi, M. Kohda, and J. Nitta, Observation of spin-splitter torque in collinear antiferromagnetic , Phys. Rev. Lett. 129, 137201 (2022).
@@
- Chi et al. [2024] B. Chi, L. Jiang, Y. Zhu, G. Yu, C. Wan, J. Zhang, and X. Han, Crystal-facet-oriented altermagnets for detecting ferromagnetic and antiferromagnetic states by giant tunneling magnetoresistance, Phys. Rev. Appl. 21, 034038 (2024).
@@
- Lander et al. [1967] G. H. Lander, P. J. Brown, and J. B. Forsyth, The antiferromagnetic structure of Mn5Si3, Proceedings of the Physical Society 91, 332 (1967).
- Menshikov et al. [1990] A. Z. Menshikov, A. P. Vokhmyanin, and Y. A. Dorofeev, Magnetic structure and phase transformations in Mn5Si3, physica status solidi (b) 158, 319 (1990).
- Brown et al. [1992] P. J. Brown, J. B. Forsyth, V. Nunez, and F. Tasset, The low-temperature antiferromagnetic structure of Mn5Si3 revised in the light of neutron polarimetry, Journal of Physics: Condensed Matter 4, 10025 (1992).
- Brown and Forsyth [1995] P. J. Brown and J. B. Forsyth, Antiferromagnetism in Mn5Si3: the magnetic structure of the AF2 phase at 70 K, Journal of Physics: Condensed Matter 7, 7619 (1995).
- Chen et al. [2021] T. Chen, T. Tomita, S. Minami, M. Fu, T. Koretsune, M. Kitatani, I. Muhammad, D. Nishio-Hamane, R. Ishii, F. Ishii, R. Arita, and S. Nakatsuji, Anomalous transport due to weyl fermions in the chiral antiferromagnets mn3x, x=sn, ge, Nature Communications 12, 572 (2021).
- Sugimoto et al. [2020] S. Sugimoto, Y. Nakatani, Y. Yamane, M. Ikhlas, K. Kondou, M. Kimata, T. Tomita, S. Nakatsuji, and Y. Otani, Electrical nucleation, displacement, and detection of antiferromagnetic domain walls in the chiral antiferromagnet mn3sn, Communications Physics 3, 111 (2020).